News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 23 September 2005 11.00am EDT

Observations were resumed on Sept 16 following last week's solar activity and the unexpected shutdown of the ACIS Digital Electronics Assembly side-A (DEA-A) on Sept 15. Initial analysis of the event indicated similarities to two prior shutdowns of the ACIS Digital Processing Assembly in 2000 and 2002 that were attributed to a cosmic ray hit to the opto-isolator resulting in a spurious command to turn off. Following a detailed engineering assessment, a real-time procedure was conducted on Sept 16 to turn the DEA-A back on, and the science program was resumed. Further analysis of the anomaly is on-going. A total of 160 Ks of science time was impacted by the solar event and DEA-A shutdown.

A real-time procedure was completed on Sept 16 to update the on-board ephemeris.

A Chandra press release was issued on 22 Sept describing observations of Tycho's supernova remnant that provide compelling evidence that a supernova shock wave has produced a large amount of cosmic rays, particles of mysterious origin that constantly bombard the Earth. For details see: http://chandra.harvard.edu/press/05_releases/press_092205.html

The schedule of targets for next week is shown below and includes a series of calibration observations of the target AR Lac.

------------------------------------------
M33 Field 5             ACIS-I       Sep 26
The Crab Nebula         ACIS-S/HETG
ELAIS-S1 A              ACIS-I
AR Lac (11 obs)         HRC-I        Sep 27
Radiation Belts
EGS-3                   ACIS-I       Sep 28
RCS2319.9+0038          ACIS-S
M33 Field 5             ACIS-I       Sep 29
EGS-3                   ACIS-I
Radiation Belts                      Sep 30
EGS-7                   ACIS-I
RCS2319.9+0038          ACIS-S       Oct 1
M33 Field 5             ACIS-I
Capella                 HRC-S/LETG   Oct 2
RCS2319.9+0038          ACIS-S
AR Lac (2 obs)          HRC-I

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports